# Minimax Decision Trees via Martingale Approximations

### v1

@@ B.2 Higher dimensional input domain
with uniformly random samples on $[0,1]^{d}$, serves as an exemplary testbed for decision tree methods in high-dimensional regression problems, offering a nuanced landscape of challenges that mirror real-world complexities. In the dense design scenario ($n>d>100$), the function’s inherent non-linearity and intricate variable interactions push the boundaries of decision trees’ capabilities. The quartic terms in the function introduce sharp curvatures in the response surface, necessitating sophisticated splitting strategies to approximate these non-linear relationships accurately. Moreover, the coupling of variables in quartets creates a web of local interactions that tests the trees’ ability to discern and model multivariable dependencies efficiently.
@@
10 4 $4.48\times 10^{2}$ $2.01\times 10^{2}$ $2.57\times 10^{2}$
10 16 $1.60\times 10^{4}$ $1.52\times 10^{4}$ $1.52\times 10^{4}$
10 64 $3.70\times 10^{4}$ $4.24\times 10^{4}$ $3.79\times 10^{4}$
10 256 $1.35\times 10^{5}$ $1.24\times 10^{5}$ $1.14\times 10^{5}$
10 1024 $9.31\times 10^{5}$ $8.04\times 10^{5}$ $4.28\times 10^{5}$
100 4 $5.49\times 10^{1}$ $5.21\times 10^{1}$ $5.31\times 10^{1}$
100 16 $2.70\times 10^{3}$ $2.67\times 10^{3}$ $2.57\times 10^{3}$
100 64 $1.99\times 10^{4}$ $1.95\times 10^{4}$ $2.20\times 10^{4}$
100 256 $1.01\times 10^{5}$ $1.02\times 10^{5}$ $9.39\times 10^{4}$
100 1024 $4.65\times 10^{5}$ $4.29\times 10^{5}$ $4.02\times 10^{5}$
1000 4 $4.68\times 10^{1}$ $4.65\times 10^{1}$ $4.49\times 10^{1}$
1000 16 $1.85\times 10^{3}$ $1.85\times 10^{3}$ $2.03\times 10^{3}$
1000 64 $1.62\times 10^{4}$ $1.62\times 10^{4}$ $1.59\times 10^{4}$
1000 256 $6.46\times 10^{4}$ $6.47\times 10^{4}$ $6.30\times 10^{4}$
1000 1024 $2.94\times 10^{5}$ $2.96\times 10^{5}$ $2.88\times 10^{5}$
10000 4 $4.61\times 10^{1}$ $4.61\times 10^{1}$ $4.63\times 10^{1}$
10000 16 $1.87\times 10^{3}$ $1.87\times 10^{3}$ $1.97\times 10^{3}$
10000 64 $1.36\times 10^{4}$ $1.36\times 10^{4}$ $1.43\times 10^{4}$
10000 256 $6.22\times 10^{4}$ $6.22\times 10^{4}$ $6.11\times 10^{4}$
10000 1024 $2.80\times 10^{5}$ $2.80\times 10^{5}$ $2.78\times 10^{5}$
Table 1: Comparison of Decision Tree Methods (max_depth=3) using the Powell function without noise. Scikit-learn’s DecisionTreeRegressor uses mean-squared Error (MSE) as its default splitting criterion, which is equivalent to minimizing the variance of the target variable within each split and serve as a baseline comparison here.
Table 1 presents a comprehensive comparison of three decision tree methods—Scikit-learn’s standard implementation, VarianceSplit and MinimaxSplit—across various sample sizes $n$ and dimensions $d$ for the Powell function. This data offers valuable insights into the performance characteristics of these methods in both dense and sparse design scenarios. In the dense design regime ($n>d$), particularly evident in the cases where the sample size (100, 1000 or 10000) exceeds the dimensionality (4, 16 or 64), we observe relatively consistent performance across all three methods. This suggests that when data is abundant relative to the input space, the choice of splitting criterion (be it the standard impurity measure used by Scikit-learn, MinimaxSplit, or VarianceSplit) has less impact on the overall performance. However, as we transition into sparse designs ($d>n$), exemplified by cases where dimensions (64, 256 or 1024) exceed the sample size (10 or 100), more pronounced differences emerge. The MinimaxSplit method often demonstrates superior performance in these high-dimensional, data-scarce scenarios. For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 8.91 × 104, outperforming both Scikit-learn (1.07 × 105) and the VarianceSplit method (9.58 × 104). This aligns with the theoretical strengths of the minimax criterion in handling sparse, high-dimensional data, where it can more robustly identify informative splits with limited samples. The VarianceSplit method generally performs comparably to Scikit-learn’s implementation, which is expected given that both likely use variants of variance reduction as their splitting criteria. However, in some sparse cases, the VarianceSplit method shows slight improvements, possibly due to specific implementation details or hyperparameter choices. As dimension increases for a fixed sample size, we observe a consistent increase in MSE across all methods, illustrating the challenges posed by the curse of dimensionality. This effect is particularly pronounced in the transition from 64 to 256 dimensions, where the MSE often increases by an order of magnitude or more. Interestingly, the performance gap between methods narrows as the sample size increases, even in high-dimensional settings. For instance, with 1000 samples and 256 dimensions, all three methods achieve comparable MSE values (around 6.4 × 104). This suggests that with sufficient data, the choice of splitting criterion becomes less critical and all methods can effectively capture the underlying structure of the Powell function.
B.3 A weighted aggregation approach
The variance approach aims to create homogeneous subsets by minimizing the average squared difference between the observed values and the mean value in each node. In contrast, the minimax criterion focuses on minimizing the maximum error within each split, potentially leading to more balanced trees. The variance criterion, similar to Scikit-learn’s MSE, seeks to minimize the overall variance in child nodes but can be implemented with different loss norms ($L^{1}$ or $L^{2}$). Cyclic minimax, a variant of minimax, alternates through features (i.e., the first and second coordinates) in a predetermined order for splitting, which can be beneficial in high-dimensional spaces when the depth is larger than the input dimension, or when feature importance is known a priori (e.g., the first coordinate is more heterogeneous).
These alternative criteria offer different trade-offs: minimax may be more robust to outliers and preserve edges better in image processing tasks, variance (especially with $L^{2}$ norm) often provides good average-case performance, and cyclic minimax can ensure a more diverse use of features. Unlike Scikit-learn’s implementation, which primarily optimizes for average error reduction, these custom criteria allow for more specialized tree structures tailored to specific problem characteristics or performance goals.
Our novel random forest (see Algorithm 1 of Appendix C) implementation diverges from conventional practices by introducing diverse splitting criteria and weighted aggregation. This approach is motivated by our previous discussion that the recognition that different splitting criteria can capture various aspects of data structure, while weighted aggregation can emphasize more accurate models in the final prediction. The implementation incorporates Variance, Minimax, and cyclic Minimax splitting methods for different weak learners within a single ensemble. The variance method optimizes for overall error reduction, minimax focuses on worst-case performance at each split, and cyclic minimax introduces a deterministic feature selection process. This diversity in the splitting criteria aims to mitigate feature importance bias present in standard random forests and enhance ensemble diversity beyond what is typically achieved through bootstrapping and random feature subset selection (Friedman, 2001).
The weighted aggregation scheme, based on the reciprocal of each tree’s training RMSE, addresses the limitation that not all trees in the ensemble are equally reliable or informative. This adaptive ensemble aggregation allows the model to adjust to the varying quality of its constituent trees, potentially leading to more robust predictions. These modifications are primarily motivated by addressing limitations of standard random forests: feature importance bias, limited ensemble diversity, and non-adaptive aggregation. By introducing varied perspectives on feature importance, adding layers of diversification, and allowing for adaptive weighting, our approach aims to create a more flexible and robust ensemble.
B.4 Heterogeneous splitting
We also develop heterogeneous splitting strategies based on our findings. The heterogeneous splitting strategy in decision trees involves using different error metrics (splitting criteria) at various depths, unlike the homogeneous strategy that uses the same metric throughout. This approach allows flexibility and adaptability, essential for handling complex datasets where different regions of the input space may benefit from different strategies. Early splits might use variance reduction for effective data division, while deeper splits could switch to minimax to minimize worst-case errors, enhancing model robustness against outliers and reducing overfitting risks; alternatively we can also alternate between VarianceSplit and MinimaxSplit to control errors (Figure 6). Additionally, strategies like cyclic MinimaxSplit leverage periodic features, potentially improving performance in time series or cyclic data. This tailored approach adapts the splitting criterion based on depth, potentially leading to more accurate and effective splits as different tree levels might require different considerations. In fact, the minimax/variance alternating strategy with $L^{2}$ norm (RMSE=0.112334) further improves the lowest RMSE compared to a homogeneous splitting criteria. Consequently, a heterogeneous strategy can lead to more accurate and robust models by optimally addressing the unique characteristics of the data at each level of the tree, making it particularly beneficial in complex datasets with variable distributions.
@@ B.4 Heterogeneous splitting
Appendix C Algorithms
Algorithm 1 Weighted Random Forest with Multiple Splitting Criteria
1:Training data $(\mathbf{X},Y)$, number of trees $N$, minimum samples to split, maximum depth, error method, loss norm, use weights flag
2:Trained Random Forest model
3:Initialize $N$ decision trees with specified parameters
4:for $i=1$ to $N$ do
5:     Set tree’s cyclic padding to $i$
6:     Create bootstrap sample $(X_{i},Y_{i})$ from $(X,Y)$
7:     Train tree using VarianceSplit, MinimaxSplit or CyclicMinimaxSplit
8:     if use_weights is True then
9:         Predict on $(X_{i},Y_{i})$ and calculate RMSE
10:         Set tree weight $w_{i}=1/\text{RMSE}$ (when we set equal weights this becomes the regular random forest)
11:     end if
12:end for
13:if use_weights is True then
14:     Normalize weights: $w_{i}=w_{i}/\sum_{j=1}^{N}w_{j}$
15:end if
16:function Predict($X_{\mathrm{test}}$)
17:     for $i=1$ to $N$ do
18:         Get predictions $p_{i}$ from tree $i$ on $X_{\mathrm{test}}$
19:     end for
20:     if use_weights = True then return $\sum_{i=1}^{N}w_{i}p_{i}$
21:     elsereturn $\frac{1}{N}\sum_{i=1}^{N}p_{i}$
22:     end if
23:end function
Algorithm 2 Single Tree fitting with different Criteria
1:function VarianceSplit(node)
2:     for each feature $j$ do
3:         Sort data points in node by feature $j$
4:         for each potential split point do
5:              Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
6:              Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
7:              Calculate total error $\leftarrow\mathrm{MSE}_{\mathrm{left}}+\mathrm{MSE}_{\mathrm{right}}$
8:         end for
9:         Find split point with minimum total error (with minimal number of splitting samples)
10:     end for
11:     return feature and split point with overall minimum error
12:end function
13:function MinimaxSplit(node)
14:     for each feature $j$ do
15:         Sort data points in node by feature $j$
16:         for each potential split point do
17:              Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
18:              Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
19:              Calculate max error $\leftarrow\max(\mathrm{MSE}_{\mathrm{left}},\mathrm{MSE}_{\mathrm{right}})$
20:         end for
21:         Find split point with minimum max error (with minimal number of splitting samples)
22:     end for
23:     return feature and split point with overall minimum max error
24:end function
25:function CyclicMinimaxSplit(node, depth, cyclic_padding)
@@ Appendix C Algorithms
27:     Sort data points in node by feature feature_index
28:     for each potential split point do
29:         Calculate $\mathrm{MSE}_{\mathrm{left}}\leftarrow\sum_{x\in\mathrm{left}}(y_{x}-\bar{y}_{\mathrm{left}})^{2}$
30:         Calculate $\mathrm{MSE}_{\mathrm{right}}\leftarrow\sum_{x\in\mathrm{right}}(y_{x}-\bar{y}_{\mathrm{right}})^{2}$
31:         Calculate max error $\leftarrow\max(\mathrm{MSE}_{\mathrm{left}},\mathrm{MSE}_{\mathrm{right}})$
32:     end for
33:     Find split point with minimum max error (with minimal number of splitting samples)
34:     return (feature_index, best split point)
35:end function

### v2

@@ D.2 Higher-dimensional input domain
with uniformly random samples on $0,1]^{d}$, serves as an exemplary testbed for decision tree methods in high-dimensional regression problems due to its inherent non-linearity and intricate variable interactions. The quartic terms in the function introduce sharp curvatures in the response surface, necessitating sophisticated splitting strategies to approximate these non-linear relationships accurately.
@@
10 4 $4.48\times 10^{2}$ $\boxed{2.01\times 10^{2}}$ $2.57\times 10^{2}$
10 16 $1.60\times 10^{4}$ $1.52\times 10^{4}$} $\boxed{1.52\times 10^{4}}$
10 64 $3.70\times 10^{4}$ $4.24\times 10^{4}$ $\boxed{3.79\times 10^{4}}$
@@
10 1024 $9.31\times 10^{5}$ $8.04\times 10^{5}$ $\boxed{4.28\times 10^{5}}$
100 4 $5.49\times 10^{1}$ $\boxed{5.21\times 10^{1}}$ $5.31\times 10^{1}$
100 16 $2.70\times 10^{3}$ $2.67\times 10^{3}$ $\boxed{2.57\times 10^{3}}$
100 64 $1.99\times 10^{4}$ $\boxed{1.95\times 10^{4}}$ $2.20\times 10^{4}$
100 256 $1.01\times 10^{5}$ $1.02\times 10^{5}$ $\boxed{9.39\times 10^{4}}$
100 1024 $4.65\times 10^{5}$ $4.29\times 10^{5}$ $\boxed{4.02\times 10^{5}}$
@@
1000 16 $\boxed{1.85\times 10^{3}}$ $1.85\times 10^{3}$ $2.03\times 10^{3}$
1000 64 $1.62\times 10^{4}$ $1.62\times 10^{4}$ $\boxed{1.59\times 10^{4}}$
1000 256 $6.46\times 10^{4}$ $6.47\times 10^{4}$ $\boxed{6.30\times 10^{4}}$
1000 1024 $2.94\times 10^{5}$ $2.96\times 10^{5}$ $\boxed{2.88\times 10^{5}}$
10000 4 $\boxed{4.61\times 10^{1}}$ $4.61\times 10^{1}$ $4.63\times 10^{1}$
10000 16 $\boxed{1.87\times 10^{3}}$ $1.87\times 10^{3}$ $1.97\times 10^{3}$
10000 64 $\boxed{1.36\times 10^{4}}$ $1.36\times 10^{4}$ $1.43\times 10^{4}$
10000 256 $6.22\times 10^{4}$ $6.22\times 10^{4}$ $\boxed{6.11\times 10^{4}}$
10000 1024 $2.80\times 10^{5}$ $2.80\times 10^{5}$ $\boxed{2.78\times 10^{5}}$
Table 1: Comparison of Decision Tree Methods ($\text{maximum depth}=3$) using the Powell function without noise. Scikit-learn’s DecisionTreeRegressor uses MSE as its default splitting criterion, which is equivalent to minimizing the variance of the target variable within each split and serves as a baseline comparison here. The winner in each setting is boxed.
Table 1 compares three decision tree methods—scikit-learn’s standard implementation, VarianceSplit and MinimaxSplit—across various sample sizes $n$ and dimensions $d$ for the Powell function. In the dense design regime ($n>d$), we observe relatively consistent performance across all three methods. However, as we transition into sparse designs ($d>n$), more pronounced differences emerge. The MinimaxSplit method often demonstrates superior performance in these high-dimensional, data-scarce scenarios. For instance, with 10 samples and 256 dimensions, the MinimaxSplit approach achieves an MSE of 8.91 × 104, outperforming both scikit-learn (1.07 × 105) and the VarianceSplit method (9.58 × 104).
Appendix E Pseudocode for algorithms
Input: Node $A_{t}$ at depth $k$ with sample $D_{t}=\{(x_{i},y_{i}):x_{i}\in A_{t}\}$; feature set $d]=\{1,\dots,d\}$ and $m_{try}=d$ by default
Output: $S_{j}(A_{t})$: candidate thresholds for feature $j$ at $A_{t}$; $A^{L}_{j,s}=A_{t}\cap\{x_{j}\leqslant s\}$, $A^{R}_{j,s}=A_{t}\cap\{x_{j}>s\}$; $P_{t}(B)=\Pr(X\in B\mid X\in A_{t})\approx n(B)/n(A_{t})$
1
VarianceSplit (CART)
  // cf. Fig. 2, left
2 for $(j,s)\in\{(j,s):j\in[d],\ s\in S_{j}(A_{t})\}$ do
3 $\mathrm{Crit}(j,s)\leftarrow P_{t}(A^{L}_{j,s})\,\mathrm{Var}(Y\mid A^{L}_{j,s})+P_{t}(A^{R}_{j,s})\,\mathrm{Var}(Y\mid A^{R}_{j,s})$
4
[…]
MinimaxSplit
  // cf. Fig. 2, middle
6 for $(j,s)\in\{(j,s):j\in[d],\ s\in S_{j}(A_{t})\}$ do
7 $\mathrm{Crit}(j,s)\leftarrow\max\big{\{}P_{t}(A^{L}_{j,s})\,\mathrm{Var}(Y\mid A^{L}_{j,s}),\ P_{t}(A^{R}_{j,s})\,\mathrm{Var}(Y\mid A^{R}_{j,s})\big{\}}$
8
[…]
Cyclic MinimaxSplit
  // cf. Fig. 2, right
$j_{k}\leftarrow 1+(k\bmod d)$
  // cyclic feature at depth $k$
10 for $s\in S_{j_{k}}(A_{t})$ do
11 $\mathrm{Crit}(j_{k},s)\leftarrow\max\big{\{}P_{t}(A^{L}_{j_{k},s})\,\mathrm{Var}(Y\mid A^{L}_{j_{k},s}),\ P_{t}(A^{R}_{j_{k},s})\,\mathrm{Var}(Y\mid A^{R}_{j_{k},s})\big{\}}$
12
$s^{\star}\leftarrow\arg\min_{s\in S_{j_{k}}(A_{t})}\mathrm{Crit}(j_{k},s)$;  split $A_{t}\to A^{L}_{j_{k},s^{\star}},\,A^{R}_{j_{k},s^{\star}}$.
Algorithm 1 Node-level splitting rules in the notation of Section 2
Figure 12: Comparison of node-level split selection in two random forest variants. Left: a random-dimension random forest selects a random feature uniformly for each node and choose the best split over the chosen feature. Right: a regular RF node samples $m_{\mathrm{try}}$ coordinates uniformly from the $d$ features and chooses the best split over all features according to the chosen impurity/criterion (the highlighted feature illustrates the winner).
Input: Sample $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{N}$, number of trees $B\in\mathbb{N}$, maximum depth $K\in\mathbb{N}$, minimum leaf size $n_{\min}\in\mathbb{N}$, per-node feature budget $m_{\mathrm{try}}$, split rule $\mathrm{SplitRule}$.
Output: Ensemble predictor $\hat{f}_{B}(x)=\frac{1}{B}\sum_{b=1}^{B}\hat{f}^{(b)}(x)$.
1
2for $b\leftarrow 1$ to $B$ do
 // Bootstrap sample
3   Draw NN indices $I^{(b)}=(i_{1},\dots,i_{N})$ uniformly with replacement from $[N]$ and set $\mathcal{D}^{(b)}=\{(X_{i_{\ell}},Y_{i_{\ell}})\}_{\ell=1}^{N}$.
 // Grow a full tree on $\mathcal{D}^{(b)}$
4   Initialize root node with index set $S_{\text{root}}=[N]$ and depth 0.
5 while there exists a node $(S,\text{depth})$ with $|S|>n_{\min}$ and $\text{depth}<K$ do
    // Sample a random subset of coordinates at this node
6      Draw without replacement a subset $\mathcal{J}\subseteq[d]$ with $|\mathcal{J}|=m_{\mathrm{try}}\in[d]$.
    // Choose best coordinate/threshold according to the chosen impurity criterion
7    $(j^{\star},t^{\star})\leftarrow\mathrm{SplitRule}(S,\mathcal{J},\mathcal{D}^{(b)})$.
    // Partition the node
8    $S_{L}\leftarrow\{i\in S:\,X_{ij^{\star}}\leqslant t^{\star}\}$, $S_{R}\leftarrow\{i\in S:\,X_{ij^{\star}}>t^{\star}\}$.
    // Create children; stop if a child is too small
9    if $|S_{L}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
10         mark left child as a leaf with prediction $\hat{\mu}_{S_{L}}=\frac{1}{|S_{L}|}\sum_{i\in S_{L}}Y_{i}$;
11    else
12         push $(S_{L},\text{depth}+1)$ to the queue;
13    if $|S_{R}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
14         mark right child as a leaf with prediction $\hat{\mu}_{S_{R}}=\frac{1}{|S_{R}|}\sum_{i\in S_{R}}Y_{i}$;
15    else
16         push $(S_{R},\text{depth}+1)$ to the queue;
17
18
 // Define the tree predictor by leaf means
19   For any $x\in\mathbb{R}^{d}$, let $\hat{f}^{(b)}(x)$ be the mean $\hat{\mu}_{S}$ of the leaf whose cell contains $x$.
Algorithm 2 Random Forest with generic split rule
Input: Sample $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{N}$, number of trees $B$, maximum depth $K$, minimum leaf size $n_{\min}$, per-node feature budget $m_{\mathrm{try}}=1$, split rule $\mathrm{SplitRule}$.
Output: Ensemble predictor $\hat{f}_{B}(x)=\frac{1}{B}\sum_{b=1}^{B}\hat{f}^{(b)}(x)$.
1
2for $b\leftarrow 1$ to $B$ do
3   Draw a bootstrap sample $\mathcal{D}^{(b)}$ of size NN from $\mathcal{D}$.
4  Initialize the root node with index set $S_{\text{root}}=[N]$ and depth 0.
5 while there exists a node $(S,\text{depth})$ with $|S|>n_{\min}$ and $\text{depth}<K$ do
    // Sample a random subset of coordinates at this node
6      Draw without replacement a subset $\mathcal{J}\subseteq[d]$ with $|\mathcal{J}|=m_{\mathrm{try}}$, where we set $m_{\mathrm{try}}=1$.
    // Choose best coordinate/threshold using only $\mathcal{J}$
7    $(j^{\star},t^{\star})\leftarrow\mathrm{SplitRule}(S,\mathcal{J},\mathcal{D}^{(b)})$.
    // Partition and create children
8    $S_{L}\leftarrow\{i\in S:\,X_{ij^{\star}}\leqslant t^{\star}\}$, $S_{R}\leftarrow\{i\in S:\,X_{ij^{\star}}>t^{\star}\}$.
9    if $|S_{L}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
10         make left child a leaf with prediction $\hat{\mu}_{S_{L}}$;
11    else
12         push $(S_{L},\text{depth}+1)$;
13    if $|S_{R}|\leqslant n_{\min}$ or $\text{depth}+1=K$ then
14         make right child a leaf with prediction $\hat{\mu}_{S_{R}}$;
15    else
16         push $(S_{R},\text{depth}+1)$;
17
18
19  Define $\hat{f}^{(b)}(x)$ by the mean of the leaf containing $x$.
Algorithm 3 Random-Dimension Random Forest (feature subsampling at each node)